For questions below refer to the following table that reflects population statistics for a hypothetical community, Jan 1 2014 to Dec 31, 2014. Show your work below.

·         Total population

·         Number of live births

·         Total deaths   

·         Number of infant deaths within the 1st year of life   

·         Number of deaths of persons aged 55 years and older

·         Number of deaths from heart disease

·         Number of deaths from cancer         

·         Number of persons with diabetes

·         200,000

·         3,500

·         1,400

·         90

·         850

·         130

·         45

·         5,600

a.      What is the age specific death rate of people aged 55 years and older?

b. What is the prevalence of diabetes?

c. What is the infant mortality rate ?

d. What is the Proportionate mortality (PMR ) for heart disease?
Answer #1

A. Age specific death rate of people aged 55years and older :

= Number of deaths in a year in specific age group / mid year population of the specific age group × 1000

Number of deaths of person aged 55years and older - 850

Total population - 200000

= 850/200000 × 1000

= 4.25%

B. Prevalence of diabetes - Number of existing cases of diabetes on a specific date ÷ Number of people in the population on this date.

= Total number of existing case of diabetes - 5600

Total population - 200000

Prevalence = 5600/200000 × 100

= 2.8%

C. Infant mortality rate - Number of deaths in a year < 1year of age / number of live births in a year × 1000

Number of infant deaths within the 1st year of life - 90

Number of live births - 3500

= 90/ 3500× 1000

= 25.71%

D. Proportionate mortality (PMR ) for heart disease :

= Total number of deaths due to heart disease / total number of deaths due to all cases × 100

Total deaths - 1400

Total number of deaths due to heart disease - 130

= 130/1400 × 100

= 9.28%